OpenStudy (anonymous):

solve: 3x=log(0.01)^10 a. -2/3 b. -1/6 c. 1/30 d 1/30,000 e. none

OpenStudy (anonymous):

\[\log(.01)=-2\] so it looks like you have \[3x=(-2)^{10}\] if i am reading it correctly

OpenStudy (anonymous):

depends on whether you mean \[\log((.01)^{10})\] or \[(\log(.01))^{10}\]

OpenStudy (anonymous):

that is, whether you mean raise .01 to the power of ten, then take the log, or take the log of .01, then raise the result to the power of ten. they are different numbers

OpenStudy (anonymous):

first one, \[.01^{10}=(10^{-2})^{10}=10^{-20}\] and so the log would be -20

OpenStudy (anonymous):

second one \[\log(.01)=-2\] and \[(-2)^{10}=1024\]

OpenStudy (anonymous):

thank you for the help. that is what i got. but i cannot figure out which multiple choice it is

OpenStudy (anonymous):

well as far as i can see there is no right answer no matter how you interpret the question

OpenStudy (anonymous):

if it is \[3x=-20\] then \[x=\frac{-20}{3}\] and it if is \[3x=1024\] then it is \[x=\frac{1024}{3}\]